OpenStudy (anonymous):

if \(A=\{2n-1:n\in\mathbb{N}\}\), \(B=\{2n:n\in\mathbb{N}\}\), \(C=\{4n:n\in\mathbb{N}\}\), \(\alpha(n)=2n\), and\[\beta(n)=\begin{cases} (n+1)/2 & \text{ if } n= odd, \\ n/2 & \text{ if } n=even, \end{cases}\]then \(\alpha(A)=\{4n-2:n\in\mathbb{N}\}\), \(\alpha(B)=\{4n:n\in\mathbb{N}\}\), and \(\alpha(C)=\{8n:n\in\mathbb{N}\}\).\[\]am i rite?

OpenStudy (anonymous):

btw with \(\mathbb{N}\) i mean \(\{1,2,...\}\)\[\]

OpenStudy (anonymous):

some people like to \(\mathbb{N}\cup\{0\}\) lol\[\]

OpenStudy (anonymous):

sure looks right to me. i take it \[\beta(n)\] is not important here

OpenStudy (anonymous):

oh yes theres part 2 lemme write it here btw thx so much

OpenStudy (anonymous):

\(\beta(A)=\beta(B)=\mathbb{N}\) and \(\beta(C)=B\)\[\]?

OpenStudy (anonymous):

mr. satellite?

OpenStudy (anonymous):

\[\beta(B)=\mathbb N\] for sure

OpenStudy (anonymous):

as is \[\beta(A)\]

OpenStudy (anonymous):

yes you have them all!

OpenStudy (anonymous):

thank you!!!